Show that if the temperature on the Celsius scale changes



by TC, the Fahrenheit temperature changes by TF -




(9/5)TC.




1
Expert's answer
2021-11-07T17:09:30-0500

We know that

Farnehite temperature and degree



centigrade temprature relationship

"T_F=\frac{9}{5}T_c+32"

Both side differenciate

"T_F=\frac{9}{5}∆T_c"

Then we can written as

"T_F=\frac{9}{5}T_c"